# Strong and noise-tolerant entanglement in dissipative optomechanics

### v3

@@ Abstract
Macroscopic entanglement, as critical quantum resources in quantum information science, has been extensively studied in coherent optomechanics over the past decades. However, entanglement in dissipative optomechanics, where the cavity linewidth is dependent on the position of the mechanical resonator, remains unexplored. In this work, we study quantum entanglement in a dissipative optomechanics, realized by a Michelson–Sagnac interferometer and a movable membrane. Such dissipative optomechanics allows to switch coherent and dissipative optomechanical coupling at will. With reliable parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power with coherent (dissipative) coupling. Further, we show that quantum entanglement generated by dissipative coupling is significantly stronger and more noise-tolerant than that generated by coherent coupling. When both the coherent and dissipative couplings are simultaneously considered, entanglement is instead weakened due to quantum interference. Our result indicates that dissipative optomechanical coupling may be a promising path to engineer strong and noise-tolerant quantum entanglement.
@@ I Introduction
Coherent optomechanics, where the cavity frequency is dependent on the position of the mechanical resonator, has garnered siginificant interest theoretically and experimentally over the past two decades Aspelmeyer et al. (2014). In such systems, various proposals have been put forward, including ground-state cooling Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), bistability Dorsel et al. (1983); Gozzini et al. (1985) and tristability Xiong et al. (2016), optomechancally induced transparency Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013), quantum squeezing Liao and Law (2011); Nunnenkamp et al. (2010); Lü et al. (2015); Aggarwal et al. (2020); Purdy et al. (2013); Luo et al. (2022), phase transition Wang et al. (2024); Jäger et al. (2019); Mann et al. (2018); Xu et al. (2017), microwave-optical conversion Guo et al. (2016); Han et al. (2021); Sahu et al. (2022); Andrews et al. (2014); Chen et al. (2023a); Xu et al. (2016), spin-optomechanical interface Xiong et al. (2021a); Chen et al. (2021); Peng et al. (2023); Xiong et al. (2023); Jäger et al. (2019), higher-order exceptional point Xiong et al. (2021b, 2022), and macroscopic quantum entanglement Vitali et al. (2007); Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023b); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b); Shang and Li (2024); Ockeloen-Korppi et al. (2018). Recently, dissipative optomechanics was theoretically proposed in a superconducting microwave circuit Elste et al. (2009), where the cavity linewidth is a function of the mechanical displacement Elste et al. (2009), giving rise to a dissipative coupling between the mechanical and cavity modes. Later, the setup was extended to a Michelson-Sagnac interferometers (MSI) with a movable membrane Xuereb et al. (2011), where coherent or dissipative optomechanical coupling is allowed to switch on and off at will by tuning the position of the membrane or the reflectivity of the beam splitter. Experimentally, dissipative optomechanics have been demonstrated in MSI Sawadsky et al. (2015); Friedrich et al. (2011), microdisc resonator coupled to a nanomechanical waveguide Li et al. (2009), photonic crystal split-beam nanocavity Wu et al. (2014), graphene drums coupled to high-Q microsphere Cole et al. (2015), and levitated sphere trapped by dual-beam optical tweezers Kuang et al. (2023). Within dissipative optomechanics, diverse intriguing quantum phenomena, such as unconventional bistability Kyriienko et al. (2014), quantum squeezing Qu and Agarwal (2015); Huang and Chen (2020), enhanced ground-state cooling Huang and Chen (2018); Liu et al. (2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al. (2009), self-sustained oscillations Huang et al. (2018), and phonon lasing Kuang et al. (2023); Zhang et al. (2022) have been investigated. However, macroscopic quantum entanglement has not been revealed yet with dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
@@
In this work, we propose to investigate quantum entanglement in an experimentally feasible dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011). The setup consists of a fixed perfect mirror and an effectively movable mirror, realized by a Michelson–Sagnac interferometer (MSI) with an integrated membrane. We begin by analyzing the steady-state mechanical displacement under both coherent and dissipative coupling. Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear dependence on the driving power, consistent with conventional optomechanics Aspelmeyer et al. (2014). In contrast, under dissipative coupling, the displacement shows a linear dependence on the driving power. Furthermore, we explore quantum entanglement in the presence of coherent and dissipative coupling. Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also remarkably more robust against bath temperature. For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5\,\text{K}$, while for entanglement with dissipative optomechanical coupling, it can be up to $\sim 25\,\text{K}$. When both the coherent and dissipative couplings are simultaneously considered, entanglement instead reduces due to quantum interference. Our work provides a promising path to generate strong and noise-resilient quantum entanglement in quantum systems with dissipative coupling.
@@ II The model and Hamiltonian
Figure 1: (a) Schematic diagram of a typical dissipative optomechanical system, realized by a perfect fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. (b) The structure of the effective movable mirror $\mathcal{M}$. The effective mirror consists of a Michelson-Sagnac interferometer (made up of a beam-splitter and two mirrors $M_{2}$ and $M_{3}$) and a movable membrane (mechanical resonator). $L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M_{1}-BS-M_{2}-M_{3}-BS}$). $a_{\rm in}$ is the input port and $a_{\rm out}$ is the output port.
As shown in Fig. 1(a), we consider an experimentally feasible dissipative optomechanics Sawadsky et al. (2015); Friedrich et al. (2011), consisting of a fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. The effective mirror can be realized by a Michelson-Sagnac interferometer (MSI) and a movable membrane between mirrors $M_{2}$ and $M_{3}$ [see Fig. 1(b)]. $L$ is the effective length of the cavity ($2L$ is the length of the Sagnac mode ${\rm M}_{1}-{\rm BS}-{\rm M}_{2}-{\rm M}_{3}-{\rm BS}$). The membrane is treated quantum mechanically, with its displacement and momentum operators denoted by $x$ and $p$, respectively, satisfying the commutation relation $[x,p]=i\hbar$. Upon the introduction of an external driving field (with frequency $\omega_{d}$ and power $P$) at the beam splitter, the system can be effectively modeled as a typical dissipative optomechanical system. Thus, both the cavity frequency $\omega_{a}(x)$ and the decay rate $\kappa_{a}(x)$ depend on the displacement $x$ of the membrane. For small membrane displacements $x$, we can Taylor expand $\omega_{a}(x)$ and $\kappa_{a}(x)$ to the first order, i.e., $\omega(x)\approx\omega_{a}-g_{\omega}x$ and $\kappa(x)\approx\kappa_{a}-g_{\kappa}x$, where $\omega_{a}$ ($\kappa_{a}$) is the frequency (decay rate) of the cavity with annihilation (creation) operator $a$ ($a^{\dagger}$) when the membrane is at its equilibrium. The Hamiltonian of the considered system is given by Elste et al. (2009); Xuereb et al. (2011)
@@
where $\omega_{m}$ is the frequency of the membrane, $\omega$ is the frequency of the bosonic mode (i.e., bath) with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$), $x^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$ is the position-dependent coupling strength between the optical cavity and the bosonic mode. In fact, the bosonic mode here is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation. At the optical domain, the input field satisfies the relations Walls and Milburn (1994)
@@
$+ix^{\prime}(A_{\rm in}a^{\dagger}-A_{\rm in}^{\dagger}a),$ (3)
@@ III Steady-state solution
$\dot{a}=$ $-(\kappa_{a}+i\Delta_{a})a+ig_{\omega}ax+g_{\kappa}ax+x^{\prime}A_{\rm in},$ (5)
where $\gamma_{m}$ is the damping rate of the membrane and $\xi$ is the Brownian noise operator arising from the coupling of the membrane oscillator to the hot environment. The correlation function of $\xi$ is given by Giovannetti and Vitali (2001) $\langle\xi(t)\xi(t^{\prime})\rangle=\frac{\gamma_{m}}{\omega_{m}}\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}\omega[{\rm coth}(\frac{\hbar\omega}{2k_{b}T})+1]$. In the limit of the high quality factor, $\omega_{m}\gg\gamma_{m}$, the Brownian noise $\xi$ becomes am approximate Markovian noise Benguria and Kac (1981), i.e.,
@@
When the evolution time is long enough, the system reaches its steady state, denoted by $a_{s}$, $x_{s}$, and $p_{s}$. From Eq. (III), the steady-state equation can be directly obtained as
@@
$\dot{a}_{s}=$ $-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E},$ (7)
where ${\Delta}_{s}=\Delta_{a}-{g_{\omega}}x_{s}$, $\kappa_{s}=\kappa_{a}-g_{\kappa}x_{s}$, and $x_{s}^{\prime}=\sqrt{2\kappa_{a}}-{g_{\kappa}x_{s}}/{\sqrt{2\kappa_{a}}}$. At the steady state, we have $\dot{\mathcal{O}}_{s}=0$, equivalently
$\omega_{m}p_{s}=$ $0,$
$g_{\omega}|a_{s}|^{2}-\omega_{m}x_{s}+\frac{ig_{\kappa}}{\sqrt{2\kappa_{a}}}(\mathcal{E}^{*}a_{s}-\mathcal{E}a_{s}^{*})-\gamma_{m}p_{s}=$ $0,$ (8)
$-(\kappa_{s}+i{\Delta}_{s})a_{s}+x_{s}^{\prime}\mathcal{E}=$ $0$
By solving these equations, we have
$p_{s}=$ $0,~{}~{}a_{s}=\frac{x_{s}^{\prime}\mathcal{E}}{\kappa_{s}+i{\Delta}_{s}},$
@@
Obivously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative coupling between photons and phonons. For a coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig. 2(a)]. For a dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
@@
where $x_{s}\neq 2\kappa_{a}/g_{\kappa}$ and $n_{s}\leq(\kappa_{a}/g_{\kappa})^{2}(\omega_{m}/2\Delta_{a})$ are both required. When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum. Here $x_{s}$ and $n_{s}$ are quadratic dependence for the dissipative coupling, demonstrated by the blue curve in Fig. 2(a). One may further notice that the displacement of the membrane caused by the coherent coupling is larger than that caused by the dissipative coupling, as clearly shown in Fig. 2(a). To plot Fig. 2(a), we adopt typical experimental parameters Sawadsky et al. (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz. These experimental parameters ensure the stability of the proposed system according to Routh-Hurwitz criterion.
@@
Explicitly, Eq. (11) may have three real roots, corresponding to bistability. But in fact, the bistability can only be predicted for coherent coupling, as shown by the red curve in Fig. 2(b). This indicates that dissipative coupling can not generate the typical bistability in optomechanics, demonstrated by the blue curve in Fig. 2(b).

### v4

@@ Abstract
Macroscopic entanglement, as a critical quantum resource in quantum information science, has been extensively studied in coherent optomechanics over the past decades. However, entanglement in dissipative optomechanics, where the cavity linewidth depends on the position of the mechanical resonator, remains largely unexplored. In this work, we investigate quantum entanglement in a dissipative optomechanical system realized by a Michelson-Sagnac interferometer with a movable membrane. This configuration enables the switching between coherent and dissipative optomechanical couplings at will. With experimentally feasible parameters, we demonstrate that the steady-state mechanical displacement exhibits a nonlinear (linear) dependence on the driving power under coherent (dissipative) coupling. Furthermore, we show that the quantum entanglement generated via dissipative coupling is significantly stronger and more robust to noise than that generated via coherent coupling. When both coherent and dissipative couplings are simultaneously present, the entanglement is weakened due to quantum interference. Our results indicate that dissipative optomechanical coupling can be a promising route for engineering strong and noise-resilient quantum entanglement.
@@ I Introduction
Coherent optomechanics, where the cavity frequency depends on the position of the mechanical resonator, has garnered significant interest both theoretically and experimentally over the past two decades Aspelmeyer et al. (2014). In such systems, various proposals have been put forward, including ground-state cooling Teufel et al. (2011); Gröblacher et al. (2009); Schliesser et al. (2008), bistability Dorsel et al. (1983); Gozzini et al. (1985) and tristability Xiong et al. (2016), optomechanically induced transparency Weis et al. (2010); Safavi-Naeini et al. (2011); Kronwald and Marquardt (2013), quantum squeezing Liao and Law (2011); Nunnenkamp et al. (2010); Lü et al. (2015); Aggarwal et al. (2020); Purdy et al. (2013); Luo et al. (2022), phase transitions Wang et al. (2024); Jäger et al. (2019); Mann et al. (2018); Xu et al. (2017), microwave-optical conversion Guo et al. (2016); Han et al. (2021); Sahu et al. (2022); Andrews et al. (2014); Chen et al. (2023a); Xu et al. (2016), spin-optomechanical interfaces Xiong et al. (2021a); Chen et al. (2021); Peng et al. (2023); Xiong et al. (2023); Jäger et al. (2019), higher-order exceptional points Xiong et al. (2021b, 2022), and macroscopic quantum entanglement Vitali et al. (2007); Tian (2012, 2013); Wang and Clerk (2013); Chen et al. (2024, 2023b); Liu et al. (2024a); Lai et al. (2022); Jiao et al. (2020, 2022); Liu et al. (2024b); Shang and Li (2024); Ockeloen-Korppi et al. (2018).
Recently, dissipative optomechanics was theoretically proposed in a superconducting microwave circuit Elste et al. (2009), where the cavity linewidth is a function of the mechanical displacement Elste et al. (2009), giving rise to a dissipative coupling between the mechanical and cavity modes. Later, the setup was extended to Michelson-Sagnac interferometers (MSIs) with a movable membrane Xuereb et al. (2011); Tarabrin et al. (2013), where coherent or dissipative optomechanical coupling can be switched on and off at will by tuning the position of the membrane or the reflectivity amplitude of the beam splitter (BS). Experimentally, dissipative optomechanics has been demonstrated in MSIs Sawadsky et al. (2015); Friedrich et al. (2011), a microdisk resonator coupled to a nanomechanical waveguide Li et al. (2009), a photonic crystal split-beam nanocavity Wu et al. (2014), graphene drums coupled to a high-Q microsphere Cole et al. (2015), and a levitated sphere trapped by dual-beam optical tweezers Kuang et al. (2023). Within dissipative optomechanics, diverse intriguing quantum phenomena, such as unconventional bistability Kyriienko et al. (2014), quantum squeezing Qu and Agarwal (2015); Huang and Chen (2020), enhanced ground-state cooling Huang and Chen (2018); Liu et al. (2023), quantum sensing Huang and Agarwal (2017), negative-damping instability Elste et al. (2009), self-sustained oscillations Huang et al. (2018), and phonon lasing Kuang et al. (2023); Zhang et al. (2022), have been investigated. However, macroscopic quantum entanglement has not yet been revealed in dissipative optomechanical coupling, although it has been intensively studied in coherent optomechanics.
@@
In this work, we investigate quantum entanglement in an experimentally realizable dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011). The setup consists of a fixed perfect mirror and an effectively movable mirror implemented using a Michelson-Sagnac interferometer (MSI) with an integrated membrane. We begin by analyzing the steady-state mechanical displacement under both coherent and dissipative coupling. Notably, with coherent coupling, the mechanical displacement exhibits a nonlinear relationship with the driving power, consistent with conventional optomechanics Aspelmeyer et al. (2014). In contrast, under dissipative coupling, the displacement shows a linear dependence on the driving power. Furthermore, we explore quantum entanglement in the presence of both coherent and dissipative coupling. Interestingly, we find that entanglement induced by dissipative coupling is not only significantly stronger than that generated by coherent coupling but also considerably more robust against bath temperature. For entanglement with coherent optomechanical coupling, its survival temperature is approximately $\sim 8.5$ K, while for entanglement with dissipative optomechanical coupling, it can reach up to $\sim 25$ K. When both coherent and dissipative couplings are simultaneously considered, entanglement is reduced due to quantum interference. Our work provides a promising path to generate strong and temperature-resilient quantum entanglement in quantum systems with dissipative coupling.
@@ II The model and Hamiltonian
Figure 1: (a) Schematic diagram of a typical dissipative optomechanical system, consisting of a perfectly fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. (b) Structure of the effective movable mirror $\mathcal{M}$. It comprises a Michelson-Sagnac interferometer, which is formed by a beam splitter and two mirrors $M_{2}$ and $M_{3}$, along with a movable membrane (mechanical resonator). The parameter $L$ denotes the effective cavity length, where $2L$ is the round-trip length of the Sagnac mode ${\rm M}_{1}$-${\rm BS}$-${\rm M}_{2}$-${\rm M}_{3}$-${\rm BS}$. The optical field enters the system through the input port $a_{\rm in}$ and exits through the output port $a_{\rm out}$.
As shown in Fig. 1(a), we consider an experimentally feasible dissipative optomechanical system Sawadsky et al. (2015); Friedrich et al. (2011), consisting of a fixed mirror $M_{1}$ and an effective movable mirror $\mathcal{M}$. The effective mirror can be realized by a Michelson-Sagnac interferometer (MSI) incorporating a movable membrane placed between mirrors $M_{2}$ and $M_{3}$ [see Fig. 1(b)]. Here, $L$ denotes the effective length of the cavity (with $2L$ representing the round-trip length of the Sagnac mode ${\rm M}_{1}$-${\rm BS}$-${\rm M}_{2}$-${\rm M}_{3}$-${\rm BS}$). The membrane is treated quantum mechanically, with its displacement and momentum operators denoted by $x$ and $p$, respectively, satisfying the canonical commutation relation $[x,p]=i\hbar$. Upon the introduction of an external driving field (with frequency $\omega_{d}$ and power $P$) at BS, the system can be effectively described as a typical dissipative optomechanical system, where both the cavity resonance frequency $\omega_{a}(x)$ and the decay rate $\kappa_{a}(x)$ depend on the membrane displacement $x$. In the good-cavity limit Yamamoto et al. (2010), the system Hamiltonian is given by Elste et al. (2009); Xuereb et al. (2011); Tarabrin et al. (2013) (setting $\hbar=1$)
@@
where $\omega_{a}$ ($\kappa_{a}$) denotes the cavity resonance frequency (decay rate) with annihilation (creation) operator $a$ ($a^{\dagger}$) when the membrane is at equilibrium, $\omega_{m}$ is the mechanical resonance frequency of the membrane, and $\omega$ is the frequency of the bosonic bath mode with annihilation (creation) operator $a_{\omega}$ ($a_{\omega}^{\dagger}$). The position-dependent coupling strength between the optical cavity and the bath is given by ${\Gamma}=\sqrt{2\kappa_{a}}-{g_{\kappa}x}/{\sqrt{2\kappa_{a}}}$. Here, $g_{\omega(\kappa)}$ represents the single-photon coherent (dissipative) optomechanical coupling strength, explicitly expressed as Elste et al. (2009); Xuereb et al. (2011); Tarabrin et al. (2013)
$g_{\omega}=$ $-2(\omega_{a}x_{\rm zpf}/L)[(|R|^{2}-|T|^{2})+\tau\cos(\arg\mathcal{T})],$
$g_{\kappa}=$ $-i\sqrt{2}(\omega_{a}x_{\rm zpf}/L)|\tau|[2RT+\rho\cos(\arg\mathcal{T})],$ (2)
with
$\rho=$ $-[(R^{2}e^{2ix}+T^{2}e^{-2ix})\mathcal{R}+2RT\mathcal{T}]e^{-i\rm{arg}\mathcal{T}},$
$\tau=$ $[(RT^{*}e^{2ix}-{\rm c.c.})\mathcal{R}-(|R|^{2}-|T|^{2})\mathcal{T}]e^{-i\rm{arg}\mathcal{T}}$ (3)
being the effective complex reflectivity and transmissivity of the mirror $\mathcal{M}$, which depend on the membrane’s position $x$. Here, $x_{\rm zpf}$ denotes the zero-point fluctuation of the membrane, and $R~{}(\mathcal{R})$ and $T~{}(\mathcal{T})$ represent the complex reflectivity and transmissivity of BS (membrane), respectively. Therefore, the values of $g_{\omega}$ and $g_{\kappa}$ can be independently tuned, allowing one to switch them on or off as desired. In particular, a purely dissipative coupling (i.e., $g_{\omega}=0$) can be achieved by ensuring $|R|\neq|T|$.
Note that the bosonic mode $a_{\omega}$ in Eq. (II) is to introduce a zero-value expectation input field $a_{\rm in}=\int\frac{d\omega}{\sqrt{2\pi}}e^{-i\omega(t-t_{0})}a_{\omega,0}$, with $a_{\omega,0}$ the operator at an initial time, to the cavity mode $a$, which can be easily obtained following the standard input-output theory with the Markov approximation. At the optical domain, the input field satisfies the relations Walls and Milburn (1994)
@@
$+i{\Gamma}(A_{\rm in}a^{\dagger}-A_{\rm in}^{\dagger}a),$ (5)
@@ III Steady-state solution
$\dot{a}=$ $-(\kappa_{a}+i\Delta_{a})a+ig_{\omega}ax+g_{\kappa}ax+{\Gamma}A_{\rm in},$ (7)
where $\gamma_{m}$ is the damping rate of the membrane and $\xi$ is the Brownian noise operator arising from the coupling of the membrane oscillator to the hot environment. Since the Brownian noise is inherently random and unbiased. In the statistical sense, its average value is zero, i.e., $\langle\xi\rangle=0$. However, investigating the fluctuation behavior of the system, the contribution from the correlation function of $\xi$ should be taken into account, given by Giovannetti and Vitali (2001) $\langle\xi(t)\xi(t^{\prime})\rangle=\frac{\gamma_{m}}{\omega_{m}}\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}\omega[{\rm coth}(\frac{\hbar\omega}{2k_{b}T})+1]$. In the limit of the high quality factor, $\omega_{m}\gg\gamma_{m}$, the Brownian noise $\xi$ approximately becomes a Markovian noise Benguria and Kac (1981), i.e.,
@@
In the long-time limit ($t\rightarrow\infty$), the system reaches a steady state. Each operator $\mathcal{O}$ is then expanded as the sum of its classical steady-state value and a small quantum fluctuation, i.e., $\mathcal{O}=\mathcal{O}_{s}+\delta\mathcal{O}$. Substituting this decomposition into Eq. (III), we derive a set of equations that govern the steady-state values $\mathcal{O}_{s}$ as follows:
@@
$\dot{a}_{s}=$ $-(\kappa_{s}+i{\Delta}_{s})a_{s}+{\Gamma_{s}}\mathcal{E},$ (9)
where ${\Delta}_{s}=\Delta_{a}-{g_{\omega}}x_{s}$, $\kappa_{s}=\kappa_{a}-g_{\kappa}x_{s}$, and ${\Gamma_{s}}=\sqrt{2\kappa_{a}}-{g_{\kappa}x_{s}}/{\sqrt{2\kappa_{a}}}$. At the steady state, we have $\dot{\mathcal{O}}_{s}=0$, yielding to
$p_{s}=$ $0,~{}~{}a_{s}=\frac{{\Gamma_{s}}\mathcal{E}}{\kappa_{s}+i{\Delta}_{s}},$
@@
Obviously, the steady-state displacement of the membrane is determined by both the dispersive and dissipative couplings between photons and phonons. For coherent coupling ($g_{\kappa}=0$), $x_{s}$ is proportional to the intracavity mean photon number $n_{s}=|a_{s}|^{2}$ [see the red curve in Fig. 2(a)]. For dissipative coupling ($g_{\omega}=0$), $x_{s}$ becomes
@@
where $x_{s}\neq 2\kappa_{a}/g_{\kappa}$ and $n_{s}\leq(\kappa_{a}/g_{\kappa})^{2}(\omega_{m}/2\Delta_{a})$ are both required. When $n_{s}=0$, $x_{s}=0$, i.e., the membrane is positioned at its equilibrum. Here $x_{s}$ exhibits a quadratic dependence on $n_{s}$ in the dissipative coupling regime, as demonstrated by the blue curve in Fig. 2(a). One can further observe that the displacement of the membrane induced by coherent coupling is larger than that induced by dissipative coupling, as clearly shown in Fig. 2(a). To plot Fig. 2(a), we employ typical experimental parameters Sawadsky et al. (2015): the wavelength of the driving field $\lambda=1064$ nm, $\omega_{d}/2\pi=c/\lambda=281.96$ THz, the quality factor of the membrane $\mathcal{Q}_{m}=5.8\times 10^{5}$, $\omega_{m}/2\pi=136$ kHz, $\gamma_{m}/2\pi=0.23$ Hz, $L=8.7$ cm, $m=80$ ng, and $\kappa_{a}/2\pi=1.5$ MHz. These parameters guarantee the stability of the proposed system according to the Routh-Hurwitz criterion.
@@
Specifically, Eq. (12) may have three real roots, corresponding to bistability. However, bistability can only be predicted for coherent coupling, as shown by the red curve in Fig. 2(b). This suggests that dissipative coupling cannot generate the typical bistability in optomechanics, as demonstrated by the blue curve in Fig. 2(b).